Find the number of atoms in a copper rod with a length of 9.20 cm and a radius of 1.00 cm. The density of copper is 8.96 g/cm3.
Answer #1

Length=9.2cm

Radius=1cm

Density=8.96g/cm^3

Vol of rod=pi× radius×radius×length=3.14×1×1×9.2=28.9cm^3

Mass=density × volume=8.96 × 28.9=259g

Molar mass of copper= 63.55 g/mol

This implies 259g will have 259/63.55 mol =4 mol

We know that 1 mol = 6.022 ×10^23 atoms

Hence 4 mol= 4×6.022 × 10^23 atoms = 2.4088 × 10^24 atoms